# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Unified Approach for Elevating Benchmark Quality

### v2

@@ 3.2. Results of the Review
CoNaLa-SO (Orlanski and Gittens, 2021) English $\rightarrow$ Python 10,000
APPS (Hendrycks et al., 2021a) English $\rightarrow$ Python 10,000
APPS-Eval (Dong et al., 2023) English $\rightarrow$ Python 10,000
AixBench (Hao et al., 2022) English, Chinese $\rightarrow$ Java 175
Natural2Code (Gemini Team, 2023) English $\rightarrow$ Python Unknown
CoSQA (Huang et al., 2021) English $\rightarrow$ Python 20,604
WebQueryTest (Lu et al., 2021) English $\rightarrow$ Python 1,046
AdvTest (Lu et al., 2021) English $\rightarrow$ Python 280,634
CONCODE (Iyer et al., 2018) English $\rightarrow$ Java 104,000
MTPB (Nijkamp et al., 2023) English $\rightarrow$ Python 115
CAASD  (Zhang et al., 2024c) English $\rightarrow$ Python 72
Shellcode_IA32  (Liguori et al., 2022) English $\rightarrow$ IA32/Shell 3200
Odex  (Liguori et al., 2022) {Spanish, Japanese, 945 {90, 164,
@@
PSB2  (Helmuth et al., 2021) English $\rightarrow$ {Clojure, Python} 25
question-answer pairs
TACO  (Li et al., 2023a) English $\rightarrow$ Python 1,539,152 on 26,433 distinct tasks
Turbulence  (Honarvar et al., 2023) English $\rightarrow$ Python 60 (with 420 total test cases)
@@
NL2ML-lib (Shin et al., 2024) English $\rightarrow$ Python (ML libraries) 11,000
RMCBench (Chen et al., 2024) English $\rightarrow$ 9 Languages 473 malicious prompts
Evil (Liguori et al., 2021) English $\rightarrow$ {Python, IA_32} 19255
Exec-CSN (Xie et al., 2024) English $\rightarrow$ Python 1,931
CodeIF (Yan et al., 2025) English $\rightarrow$ {Java, Python, Go, C++} 1,200 tasks
CodeIF-Bench (Wang et al., 2025e) English $\rightarrow$ Python 122 tasks (42 repositories)
ARCADE (Yin et al., 2023) English $\rightarrow$ Python 1082
StackEval (Shah et al., 2024) English $\rightarrow$ Python 925
SwiftEval (Petrukha et al., 2025) English $\rightarrow$ Swift 28
CoSQA+ (Gong et al., 2025) English $\rightarrow$ Python 412,080 pairs
Table 7. Overview of Natural Language Benchmarks (Continued).
Category Name Language(s) No. of Problems
Text2Text (about code) InfiCoder-Eval (InfiCoder, 2023) English $\rightarrow$ English 270
BRIGHT (Su et al., 2024) English $\rightarrow$ English 1,398
Code2Text DeepCom (Hu et al., 2018) Java $\rightarrow$ English 588K
Hybrid-DeepCom (Hu et al., 2020) Java $\rightarrow$ English 466k
BinSum (Jin et al., 2023) Binary functions $\rightarrow$ English 557K
Code Attention  (Allamanis et al., 2016) Java $\rightarrow$ English 11 projects
Funcom  (LeClair et al., 2019) Java $\rightarrow$ English 2.1M problems
CodeSum (Hu et al., [n. d.]) Java $\rightarrow$ English 410,630
CoDesc (Hasan et al., 2021) Java $\rightarrow$ English 4.21M datapoints
Parallel (Barone and Sennrich, [n. d.]) Python $\rightarrow$ English 150k function/doc pais
CoDocBench (Pai et al., 2025) Python $\rightarrow$ English 4573 code/doc pairs
PoorCodeSum (Hu et al., 2024) {Java, Python, Go} {10,955 , 14,918 , 8,122}
@@
P-CodeSum (Yun et al., 2024) {Python, Java, Go, 1,500 pairs
JS, PHP, Ruby} $\rightarrow$ English
Table 8. Overview of SQL-related Benchmarks.
Category Name Language(s) No. of Problems
Text2Code BIRD  (Li et al., 2023b) English $\rightarrow$ SQL 12,751
KaggleDBQA  (Lee et al., 2021) English $\rightarrow$ SQL 272, paired with golden solutions
StacQc  (Yao et al., 2018) English $\rightarrow$ {Python/SQL} {147,546 / 119,519}
question-answer pairs
Spider(V2111111see Table 13(Lei et al., 2024) English $\rightarrow$ SQL 632 queries
Spider-Syn (Gan et al., 2021b) English $\rightarrow$ SQL (7000 / 1034)
Spider-Real (Deng et al., 2021) English $\rightarrow$ SQL 508
Spider-DK (Gan et al., 2021a) English $\rightarrow$ SQL 535 pairs
Spider-CN (Min et al., 2019) Chinese $\rightarrow$ SQL 9691 queries
SParC (Yu et al., 2019b) English $\rightarrow$ SQL 4,298 question sequences
Lyra (Liang et al., 2022) {English, Chinese} $\rightarrow$ {python, SQL} 2000
DuSQL (Wang et al., 2020) Chinese $\rightarrow$ SQL 23,797 question/SQL pairs
CoSQL (Yu et al., 2019a) English $\rightarrow$ SQL 3,007 Question Sequences
SynSQL-2.5M (Li et al., 2025a) English $\rightarrow$ SQL 2,544,390
PAUQ (Bakshandaeva et al., 2022) Russian $\rightarrow$ SQL 9,691
Ar-Spider (Almohaimeed et al., 2024) Arabic $\rightarrow$ SQL 9,691
Tur2SQL (Kanburoğlu and Tek, 2023) Turkish $\rightarrow$ SQL 10,809 question/SQL pairs
@@
Table 3 features benchmarks with competitive programming as their root, i.e., those used for understanding code complexity and efficiency. Table 4 features a set of benchmarks specifically designed to evaluate the performance of models on Data Science-related tasks along with some other domain-specific SE tasks. A notable example in this table is the Bio-Coder series of benchmarks specifically designed for bioinformatics tasks. To assess the mathematical reasoning capabilities of AI4SE models, see Table 5. Besides numbers and code, natural language is also a key component in AI4SE. From supporting instruction-tuned AI4SE models, which align more with the human brain (Aw et al., 2023), that aim to accomplish question and answering (QA) similar to the widely recognized platform StackOverflow, to summarizing code and generating tags, Table 6 and Table 7 feature AI4SE benchmarks including natural language: text-to-code, code-to-text and text-to-text (code related). We have intentionally isolated all SQL-related benchmarks due to their abundance and to facilitate locating the correct benchmark; these are presented in Table 8.
@@
Table 9. Overview of Selected Real-to-Life SE Benchmarks. (Note: X/Y/Z denotes Train/Dev/Test)
Category Benchmark Language(s) No. of Problems
@@
Software Development & Agent Benchmarks DevBench (Li et al., 2024e) Python, C/C++, Java, JavaScript 22 repositories
DevEval (Li et al., 2024d) Python 1,874
CoderUJB (Zeng et al., 2024a) Java 2,239
CODAL (Weyssow et al., 2024) Python 500
ToolQA (Zhuang et al., 2023) Python, Math, English 800(Easy)/730(Hard)
MIT (Wang et al., 2024d) Python, English 586 Problems
SAFIM (Gong et al., 2024) Python, Java, C++, C# 17720
AgentBench (Liu et al., 2023d) N/A 1360 prompts
CSR-Bench (Xiao et al., 2025) Python 100 repositories
Class Level ClassEval (Du et al., 2023) Python 100
CONCODE (Iyer et al., 2018) English, Java 104,000
BigCodeBench (Zhuo et al., 2024) Python 1,140
OOP-Bench (Wang et al., 2024b) Python 431
CodeSense (Roy et al., 2025) Python, C, Java 2125 (Python), 876 (C)
@@
ClassEval-T (Xue et al., 2025) Java, C++ 1,243)
Project & Cross-file SWE-bench (Jimenez et al., 2023) Python 19,008 (Train), 225 (Dev),
@@
CrossCodeEval (Ding et al., 2023) C#, TypeScript, Java, Python 2,665 (Python), 2,139 (Java),
3,356 (TypeScript), 1,768 (C#)
CoderEval (Yu et al., 2023) Java, Python 230
DotPrompts (Agrawal et al., 2023) Java 105538 problems
@@
BigCloneBench (Svajlenko and Roy, 2015) Java 25,000 Java Systms
DI-Bench (Zhang et al., 2025c) Python, C#, Rust, JS 581 repositories
@@
DyPyBench (Bouzenia et al., 2024) Python 50 repositories
Multi-SWE Java, TypeScript, JavaScript 500 (Python), 128 (Java),
-bench (Zan et al., 2025) Go, Rust, C, C++ 224 (TS), 356 (JS),
@@
KernelBench (Ouyang et al., 2025) Python 250
CodeMEnv (Cheng et al., 2025) Python, Java 587 (Python), 335 (Java)
CodeEditorBench (Guo et al., 2025) Python, Java, C++ 7,961
ProjectEval (Liu et al., 2025a) Python 284
Table 10. Overview of Selected Real-to-Life SE Benchmarks. (Continued - Note: X/Y/Z denotes Train/Dev/Test)
Category Benchmark Language(s) No. of Problems
Repository Level RepoBench (Liu et al., 2023b) Python, Java Cross-file: 8,033, In-file: 7,910
RepoEval (Zhang et al., 2023a) Python 1,600 (line),
1,600 (API), 373 (function)
EvoCodeBench (Li et al., 2024c) Python 275
SketchEval (Zan et al., 2024) Python 19 repositories
(5 easy, 8 medium, 6 hard)
Stack Repo (Zan et al., 2024) Python (435,890 / 220,615 / 159,822)
answer pairs
ML-BENCH (Tang et al., 2024a) Python & Bash 9641 problems
CodeGen4Libs (Liu et al., 2023c) Java 403,780 prompts
SWE-rebench (Badertdinov et al., 2025) Python 294 tasks, 169 repositories
SWE-Polybench (Rashid et al., 2025) Java, JavaScript 165 (Java), 1017 (JavaScript),
TypeScript, Python 729 (TypeScript), 199 (Python)
HumanEvo (Zheng et al., 2025b) Python, Java 200 (Python), 200 (Java)
REPOCOD (Liang et al., 2024) Python 980 functions (11 projects)
FEA-Bench (Li et al., 2025b) Python 83 repositories
JavaBench (Cao et al., 2024) Java 389 methods (106 classes)
SWE-bench live (Zhang et al., 2025b) Java 1,319 tasks (93 repositories)
SWE-Lancer (Miserendino et al., 2025) N/A 1,488 tasks)
Table 11. Overview of Selected API and Retrieval Benchmarks by Category.
Category Benchmark Sources/API(s) No. of Problems
API Prediction RestBench (Song et al., 2023) Spotify, TMDB 57, 100
APIBench-Q (Peng et al., 2021) StackOverflow, Tutorial Websites 6,563 (Java),
@@
BIKER (Huang et al., 2018) StackOverflow 33,000
Gorilla APIBench (Patil et al., 2023) HuggingFace, TensorHub, TorchHub 925, 696, 94
Gorilla APIZoo (Patil et al., 2023) Open submissions
@@
Retrieval & Planning API-Bank (Li et al., 2023d) 73 commonly used APIs 753
CodeRAG-Bench (Wang et al., 2024a) Competition solutions, tutorials, 25,859
@@
Search4Code (Rao et al., 2021) Bing 6596(java)/4974(c#)
CoIR (Li et al., 2024a) GitHub, StackOverflow, and 2.38M (corpus)
@@
Memorization SATML-ext (Al-Kaswan et al., 2024) GitHub 1,000 samples
API Misuse Detection ExampleCheck (Zhang et al., 2018) StackOverflow 100 (Java)
ROBUSTAPI (Zhong and Wang, 2024) StackOverflow 1208 (18 Java APIs)
APIMU4C (Gu et al., 2019) Juliet Test Suite, ITC, 2272 (C)
@@
Table 12. Overview of AI4SE Benchmarks Related to Pseudocode.
Category Benchmark Language(s) No. of Problems Crowdsourced
Pseudocode to Code SPoC (Kulal et al., 2019) C++ 18,356 Yes
NAPS (Zavershynskyi et al., 2018) Java/UAST 17,477 No
PseudoEval (Wu et al., 2025) Python, C++, Rust 1,060 No
Code to Pseudocode Django (Oda et al., 2015) Python, English 18,805 (Train), 1,000 (Dev), No
@@
Table 13. Overview of Selected Crowd-sourced Benchmarks.
Benchmark Language(s) No. of Problems Source
WikiSQL (Zhong et al., 2017) Natural language $\rightarrow$ SQL query 80,654 Amazon MTurk (deprecated - 2017)
Spider (Yu et al., 2018) Natural language $\rightarrow$ SQL query 10,181 11 Yale students (2018)
NL2Bash (Lin et al., 2018) Natural language $\rightarrow$ Bash 9,305 Upwork (2018)
NAPS (Zavershynskyi et al., 2018) Java/UAST $\rightarrow$ Pseudocode 17,477 Self-hosted crowdsourcing,
comp. programming community (2018)
SPoC (Kulal et al., 2019) C++ 18,356 comp. programming websites (2019)
MBPP (Austin et al., 2021) Python 974 Google Research,
@@
While translating natural language is more trivial nowadays, translating code remains challenging due to various reasons (e.g. versioning, semantics, dependencies). With the lack of diversity in language support for AI4SE benchmarks and also benefiting numerous other SE tasks, Table 14 features an overview of resources that can support the ongoing development of code translation.
Table 14. Overview of Programming Language Translation Benchmarks (Note: X/Y/Z denotes Train/Dev/Test).
Category Name Language(s) No. of Samples
Programming Languages CodeTrans (Lu et al., 2021) C#, Java 11,800
TransCoder-ST (Roziere et al., 2022) C++, Java, Python 437,030
CoST (Zhu et al., 2022b) 7 programming languages 16,738
AVATAR (Ahmad et al., 2023) Java, Python 7,133 / 476 / 1,906
Multilingual-Trans (Yan et al., 2023) 8 programming languages 30,419 total
NicheTrans (Yan et al., 2023) Various niche languages 236,468 total
LLMTrans (Yan et al., 2023) 8 programming languages 350
G-TransEva (Jiao et al., 2023) 5 programming languages 400 total
CODEDITOR (Zhang et al., 2023b) C# & Java 6613
RustRepoTrans (Ou et al., 2024) C++, Java, Python $\rightarrow$ Rust 375
AVATAR-TC (Jana et al., 2024) Java, Python 55,179 / 443 / 1,746
RustRepoTrans (Ou et al., 2024) C++, Java, Python $\rightarrow$ Rust 375
RepoTransBench (Wang et al., 2024e) Python $\rightarrow$ Java 100 repositories
Libraries DLTrans (Yan et al., 2023) PyTorch, TensorFlow,
@@
Intermediate Representation SLTrans (Paul et al., 2024) 14 Languages $\rightarrow$ LLVM-IR 4M
Language Conversion Frameworks MultiPL-E (Cassano et al., 2022) 19 programming languages -
MultiEval (Athiwaratkun et al., 2022) 13 programming languages -
However, the workflow of a developer is not merely an exercise of writing snippets of code for given descriptions, but rather having a general overview of a project and how one can implement functionality such that it fits well into a collective code base. For this, Table 9 and Table 10 provide a collection of benchmarks that examine the capabilities of models to generate code on a larger scale.
The utilization of APIs plays a significant role in AI4SE benchmarks, specifically for models with Retrieval Augmented Generation (RAG) capabilities. In Table 11, prominent benchmarks focusing on leveraging the power of APIs are denoted. Furthermore, Table 12 lists benchmarks related to pseudocode, followed by an overview of notable crowd-sourced AI4SE resources in Table 13.
Table 15. Overview of Automated Program Repair, Fault Localization, and Vulnerability Detection Benchmarks.
@@
Automated Program Repair & Fault Localization Defects4J (Just et al., 2014) Java 835
GitBug-Java (Silva et al., 2024) Java 199
EvalGPTFix  (Zhang et al., 2023c) Java 4530
TutorCode  (Yang et al., 2024) C++ 1239
GHRB  (Lee et al., 2023) Java 107
IntroClass  (Goues et al., 2015) C 998
ManyBugs  (Goues et al., 2015) 7 Languages 185
DebugBench (Tian et al., 2024) C++, Java, Python 1,438 & 1,401 & 1,414
QuixBugs (Lin et al., 2017) Java 40 (locations of bugs)
RES-Q (Labash et al., 2024) Python, JS 100 hand-crafted questions
+ test scripts
StudentEval (Babe et al., 2023) Python 1,749 buggy programs (48 tasks)
(3 test cases per problem)
Re-Factory (Hu et al., 2019) Python 1783(buggy)/2442(correct)
ConDefects (Wu et al., 2023b) Python, Java 526(Python), Java(477)
Cerberus (Shariffdeen et al., 2023) C, C++, Java 2242 (across 4 task types)
RepairBench (Silva and Monperrus, 2025) Java 574
MaRV (Nunes et al., 2025) Java 693
BugsinPy (Widyasari et al., 2020) Python 493 bugs
Vulnerability Detection CVEFixes (Bhandari et al., 2021) Various Languages 5,365
LLMSecEval (Tony et al., 2023) C 150 (on 25 distinct vulnerabilities)
SecurityEval (Siddiq et al., 2022) 6 languages 130 (on 75 common vulnerabilities)
Vul4J (Bui et al., 2022) Java 79 vulnerabilities
FormAI (Tihanyi et al., 2023) C 112,000 labeled instances
VJBbench (Wu et al., 2023a) Java 42 vulnerabilities
SmartBugs (Durieux et al., 2020) Solidity 69 Vulnerable Smart Contracts
Devign (Zhou et al., 2019) C 4 large-scale software repositories
D2A (Zheng et al., 2021) C/C++ 6 OSS Programs
BigVul (Fan et al., 2020) C/C++ 348 Projects
SARD  [^12] https://samate.nist.gov/SARD/ Java, C, C++, C#, PHP 32k [^13] As of 4th Feb 2025
NVD  [^16] https://nvd.nist.gov/developers/data-sources Various Languages 265k [^17] As of 4th Feb 2025
ARVO (Mei et al., 2024) C, C++ 1,001 vulnerabilities
VADER (Liu et al., 2025b) 15 languages 174 vulnerabilities
ManyVuls4J (Lin et al., 2024b) Java 103 vulnerabilities
Software Testing CoverageEval (Tufano et al., 2023) Python 1160
ATLAS (Watson et al., 2020) Java 9,275 projects
HITS (Wang et al., 2024c) Java 10 projects
MeMo (Bareiß et al., 2022) Java 9 projects
MLAPIs (Wan et al., 2022) Python 63 applications
CoderUJB (Zeng et al., 2024b) Java 2,239
TestBench (Zhang et al., 2024b) Java 108
TestEval (Wang et al., 2024f) Python 210
TARBENCH (Yaraghi et al., 2025) Java 45,373 (59 projects)
ProjectTest (Wang et al., 2025c) Python, Java, Javascript 20 per language
CLOVER (Xu et al., 2025c) Python 845
Table 16. Overview of Selected SE-Workflow Benchmarks.
Category Benchmark Language(s) No. of Samples
Code Synthesis & Understanding Methods2Test (Tufano et al., 2020) Java 780,944
CRUXEval (Gu et al., 2024) Python 800
CRQBench (Dinella et al., 2024) C++ 100
CriticBench (Lin et al., 2024a) Python 3,825(across 5 tasks)
CodeScope (Yan et al., 2024) 8 PLs 13,390 (across 8 tasks)
CodeCriticBench (Zhang et al., 2025a) Various Languages 1,517 (Easy), 1,084 (Medium),

### v3

@@ III-B Results of the Review
CoNaLa-SO [54] English $\rightarrow$ Python 10,000
APPS [17] English $\rightarrow$ Python 10,000
APPS-Eval [9] English $\rightarrow$ Python 10,000
AixBench [55] English, Chinese 175
$\rightarrow$ Java
Natural2Code [2] English $\rightarrow$ Python Unknown
CoSQA [56] English $\rightarrow$ Python 20,604
WebQueryTest [57] English $\rightarrow$ Python 1,046
AdvTest [57] English $\rightarrow$ Python 280,634
CONCODE [58] English $\rightarrow$ Java 104,000
MTPB [59] English $\rightarrow$ Python 115
CAASD  [60] English $\rightarrow$ Python 72
Shellcode_IA32  [61] English $\rightarrow$ IA32/Shell 3200
Odex  [61] {Spanish, Japanese, 945 {90, 164,
@@
PSB2  [62] English $\rightarrow$ 25
{Clojure, Python} question-answer pairs
TACO  [63] English $\rightarrow$ Python 1,539,152 on 26,433
distinct tasks
Turbulence  [64] English $\rightarrow$ Python 60 (with 420
total test cases)
@@
NL2ML-lib [65] English $\rightarrow$ Python 11,000
RMCBench [66] English 473 malicious
$\rightarrow$ 9 Languages prompts
Evil [67] English 19255
$\rightarrow$ {Python, IA_32}
Exec-CSN [68] English $\rightarrow$ Python 1,931
CodeIF [69] English $\rightarrow$ 1,200 tasks
{Java, Python,
Go, C++}
CodeIF-Bench [70] English $\rightarrow$ Python 122 tasks
(42 repositories)
ARCADE [71] English $\rightarrow$ Python 1082
StackEval [72] English $\rightarrow$ Python 925
SwiftEval [73] English $\rightarrow$ Swift 28
CoSQA+ [74] English $\rightarrow$ Python 412,080 pairs
TABLE VII: Overview of Natural Language Benchmarks (Continued). The First two benchmarks are Text2Text (related to code) benchmarks, while the rest are Code2Text benchmarks.
Name Language(s) # Problems
InfiCoder-Eval [75] English $\rightarrow$ English 270
BRIGHT [51] English $\rightarrow$ English 1,398
DeepCom [76] Java $\rightarrow$ English 588K
Hybrid-DeepCom [77] Java $\rightarrow$ English 466k
BinSum [78] Binary functions 557K
@@
Code Attention  [79] Java $\rightarrow$ English 11 projects
Funcom  [80] Java $\rightarrow$ English 2.1M problems
CodeSum [81] Java $\rightarrow$ English 410,630
CoDesc [82] Java $\rightarrow$ English 4.21M datapoints
Parallel [83] Python $\rightarrow$ English 150k function/doc pais
CoDocBench [84] Python $\rightarrow$ English 4573 code/doc pairs
PoorCodeSum [85] {Java, Python, Go} {10,955 , 14,918
$\rightarrow$ English , 8,122}
P-CodeSum [86] {Python, Java, Go, 1,500 pairs
JS, PHP, Ruby}
$\rightarrow$ English
TABLE VIII: Overview of SQL-related Benchmarks.
Name Language(s) # Problems
BIRD  [87] English $\rightarrow$ SQL 12,751
KaggleDBQA  [88] English 272, paired with
$\rightarrow$ SQL golden solutions
StacQc  [89] English {147,546 / 119,519}
$\rightarrow$ {Python/SQL} question-answer pairs
Spider(V2111111see Table XIII[90] English $\rightarrow$ SQL 632 queries
Spider-Syn [91] English $\rightarrow$ SQL (7000 / 1034)
Spider-Real [92] English $\rightarrow$ SQL 508
Spider-DK [93] English $\rightarrow$ SQL 535 pairs
Spider-CN [94] Chinese $\rightarrow$ SQL 9691 queries
SParC [95] English $\rightarrow$ SQL 4,298 question sequences
Lyra [96] {English, Chinese} 2000
$\rightarrow$ {python, SQL}
DuSQL [97] Chinese 23,797
$\rightarrow$ SQL question/SQL pairs
CoSQL [98] English $\rightarrow$ SQL 3,007 Question Sequences
SynSQL-2.5M [99] English $\rightarrow$ SQL 2,544,390
PAUQ [100] Russian $\rightarrow$ SQL 9,691
Ar-Spider [101] Arabic $\rightarrow$ SQL 9,691
Tur2SQL [102] Turkish $\rightarrow$ SQL 10,809 question/SQL pairs
@@
Table III features benchmarks with competitive programming as their root, i.e., those used for understanding code complexity and efficiency. Table IV features a set of benchmarks specifically designed to evaluate the performance of models on Data Science-related tasks along with some other domain-specific SE tasks. A notable example in this table is the Bio-Coder series of benchmarks specifically designed for bioinformatics tasks. To assess the mathematical reasoning capabilities of AI4SE models, see Table V. Besides numbers and code, natural language is also a key component in AI4SE. From supporting instruction-tuned AI4SE models, which align more with the human brain [103], that aim to accomplish question and answering (QA) similar to the widely recognized platform StackOverflow, to summarizing code and generating tags, Table VI and Table VII feature AI4SE benchmarks including natural language: text-to-code, code-to-text and text-to-text (code related). We have intentionally isolated all SQL-related benchmarks due to their abundance and to facilitate locating the correct benchmark; presented in Table VIII.
@@
TABLE IX: Overview of Selected Real-to-Life SE Benchmarks. (Note: X/Y/Z denotes Train/Dev/Test)
Category Benchmark Language(s) # Problems
@@
Software Development & Agent Benchmarks DevBench [104] Python, C/C++, Java, JS 22 repositories
DevEval [105] Python 1,874
CoderUJB [106] Java 2,239
CODAL [31] Python 500
ToolQA [107] Python, Math, English 800(Easy)/730(Hard)
MIT [108] Python, English 586 Problems
SAFIM [109] Python, Java, C++, C# 17720
AgentBench [110] N/A 1360 prompts
CSR-Bench [111] Python 100 repositories
Class Level ClassEval [112] Python 100
CONCODE [58] English, Java 104,000
BigCodeBench [113] Python 1,140
OOP-Bench [114] Python 431
CodeSense [115] Python, C, Java 2125 (Python), 876 (C)
@@
ClassEval-T [116] Java, C++ 1,243)
Project & Cross-file SWE-bench [117] Python 19,008 (Train), 225 (Dev),
@@
CrossCodeEval [118] C#, TS, Java, Python 2,665 (Python), 2,139 (Java),
3,356 (TS), 1,768 (C#)
CoderEval [119] Java, Python 230
DotPrompts [120] Java 105538 problems
@@
BigCloneBench [121] Java 25,000 Java Systms
DI-Bench [122] Python, C#, Rust, JS 581 repositories
@@
DyPyBench [123] Python 50 repositories
Multi-SWE Java, TS, JS 500 (Python), 128 (Java),
-bench [124] Go, Rust, C, C++ 224 (TS), 356 (JS),
@@
KernelBench [125] Python 250
CodeMEnv [126] Python, Java 587 (Python), 335 (Java)
CodeEditorBench [127] Python, Java, C++ 7,961
ProjectEval [128] Python 284
TABLE X: Overview of Selected Real-to-Life SE Benchmarks. This table contains the Repository-Level benchmarks. (Continued - Note: X/Y/Z denotes Train/Dev/Test)
Benchmark Language(s) # Problems
RepoBench [129] Python, Java Cross-file: 8,033
In-file: 7,910
RepoEval [130] Python 1,600 (line),
1,600 (API)
373 (function)
EvoCodeBench [131] Python 275
SketchEval [132] Python 19 repositories
(5 easy, 8 medium
6 hard)
Stack Repo [132] Python (435,890 / 220,615
/ 159,822) answer pairs
ML-BENCH [133] Python & Bash 9641 problems
CodeGen4Libs [134] Java 403,780 prompts
SWE-rebench [135] Python 294 tasks, 169 repositories
SWE-Polybench [136] Java 165 (Java)
JS 1017 (JS),
TS 729 (TS)
Python 729 (199 (Python)
HumanEvo [137] Python, Java 200 (Python), 200 (Java)
REPOCOD [138] Python 980 functions (11 projects)
FEA-Bench [139] Python 83 repositories
JavaBench [140] Java 389 methods (106 classes)
SWE-bench live [141] Java 1,319 tasks (93 repositories)
SWE-Lancer [142] N/A 1,488 tasks)
TABLE XI: Overview of Selected API and Retrieval Benchmarks by Category.
Category Benchmark Sources/API(s) # Problems
API Prediction RestBench [143] Spotify, TMDB 57, 100
APIBench-Q [144] StackOverflow, Tutorial Websites 6,563 (Java),
@@
BIKER [145] StackOverflow 33,000
Gorilla APIBench [146] HuggingFace, TensorHub, TorchHub 925, 696, 94
Gorilla APIZoo [146] Open submissions
@@
Retrieval & Planning API-Bank [147] 73 commonly used APIs 753
CodeRAG-Bench [148] Competition solutions, tutorials, 25,859
@@
Search4Code [149] Bing 6596(java)/4974(c#)
CoIR [150] GitHub, StackOverflow, and 2.38M (corpus)
@@
Memorization SATML-ext [151] GitHub 1,000 samples
API Misuse Detection ExampleCheck [152] StackOverflow 100 (Java)
ROBUSTAPI [153] StackOverflow 1208 (18 Java APIs)
APIMU4C [154] Juliet Test Suite, ITC, 2272 (C)
@@
TABLE XII: Overview of AI4SE Benchmarks Related to Pseudocode.
Category Benchmark Language(s) # Problems Crowdsourced
Pseudocode to Code SPoC [155] C++ 18,356 Yes
NAPS [156] Java/UAST 17,477 No
PseudoEval [157] Python, C++, Rust 1,060 No
Code to Pseudocode Django [158] Python, English 18,805 (Train), 1,000 (Dev), No
@@
TABLE XIII: Overview of Selected Crowd-sourced Benchmarks (NL denotes Natural Language).
Benchmark Language(s) # Problems Source
[…]
programming community (2018)
SPoC [155] C++ 18,356 programming websites (2019)
MBPP [15] Python 974 Google Research,
@@
While translating natural language is more trivial nowadays, translating code remains challenging due to various reasons (e.g. versioning, semantics, dependencies). With the lack of diversity in language support for AI4SE benchmarks and also benefiting numerous other SE tasks, Table XIV features an overview of resources that can support the ongoing development of code translation.
TABLE XIV: Overview of Programming Language Translation Benchmarks (Note: X/Y/Z denotes Train/Dev/Test).
Category Name Language(s) # Samples
Programming Languages CodeTrans [57] C#, Java 11,800
TransCoder-ST [162] C++, Java, Python 437,030
CoST [163] 7 programming languages 16,738
AVATAR [164] Java, Python 7,133 / 476 / 1,906
Multilingual-Trans [165] 8 programming languages 30,419 total
NicheTrans [165] Various niche languages 236,468 total
LLMTrans [165] 8 programming languages 350
G-TransEva [166] 5 programming languages 400 total
CODEDITOR [167] C# & Java 6613
RustRepoTrans [168] C++, Java, Python $\rightarrow$ Rust 375
AVATAR-TC [169] Java, Python 55,179 / 443 / 1,746
RustRepoTrans [168] C++, Java, Python $\rightarrow$ Rust 375
RepoTransBench [170] Python $\rightarrow$ Java 100 repositories
Libraries DLTrans [165] PyTorch, TensorFlow,
@@
Intermediate Representation SLTrans [171] 14 Languages $\rightarrow$ LLVM-IR 4M
Language Conversion Frameworks MultiPL-E [6] 19 programming languages -
MultiEval [5] 13 programming languages -
However, the workflow of a developer is not merely an exercise of writing snippets of code for given descriptions, but rather having a general overview of a project and how one can implement functionality such that it fits well into a collective code base. For this, Table IX and Table X provide a collection of benchmarks that examine the capabilities of models to generate code on a larger scale.
The utilization of APIs plays a significant role in AI4SE benchmarks, specifically for models with Retrieval Augmented Generation (RAG) capabilities. In Table XI, prominent benchmarks focusing on leveraging the power of APIs are denoted. Furthermore, Table XII lists benchmarks related to pseudocode, followed by an overview of notable crowd-sourced AI4SE resources in Table XIII.
TABLE XV: Overview of Automated Program Repair, Fault Localization, and Vulnerability Detection Benchmarks.
Benchmark Language(s) #Samples
Automated Program Repair & Fault Localization Defects4J [172] Java 835
GitBug-Java [173] Java 199
EvalGPTFix  [174] Java 4530
TutorCode  [175] C++ 1239
GHRB  [176] Java 107
IntroClass  [177] C 998
ManyBugs  [177] 7 Languages 185
DebugBench [178] C++, Java 1,438 & 1,401
Python & 1,414
QuixBugs [179] Java 40 (locations of bugs)
RES-Q [180] Python, JS 100 hand-crafted
questions + tests
StudentEval [181] Python 1,749 buggy programs
(48 * 3 tests)
Re-Factory [182] Python 1783(buggy)
2442(correct)
ConDefects [183] Python 526(Python)
Java Java(477)
Cerberus [184] C, C++, Java 2242 (across 4 tasks)
RepairBench [185] Java 574
MaRV [186] Java 693
BugsinPy [187] Python 493 bugs
Vulnerability Detection CVEFixes [188] Various 5,365
LLMSecEval [189] C 150 (on 25
vulnerabilities)
SecurityEval [190] 6 languages 130 (on 75
vulnerabilities)
Vul4J [191] Java 79 vulnerabilities
FormAI [192] C 112k instances
VJBbench [193] Java 42 vulnerabilities
SmartBugs [194] Solidity 69 Vulnerable
Smart Contracts
Devign [195] C 4 large
Software Repos
D2A [196] C/C++ 6 OSS Programs
BigVul [197] C/C++ 348 Projects
SARD  [^12] https://samate.nist.gov/SARD/ Java, C, C++ 32k [^13] As of 4th Feb 2025
C#, PHP
NVD  [^16] https://nvd.nist.gov/developers/data-sources Various 265k [^17] As of 4th Feb 2025
ARVO [198] C, C++ 1,001 vulnerabilities
VADER [199] 15 languages 174 vulnerabilities
ManyVuls4J [200] Java 103 vulnerabilities
Software Testing CoverageEval [201] Python 1160
ATLAS [202] Java 9,275 projects
HITS [203] Java 10 projects
MeMo [204] Java 9 projects
MLAPIs [205] Python 63 applications
CoderUJB [206] Java 2,239
TestBench [207] Java 108
TestEval [208] Python 210
TARBENCH [209] Java 45,373 (59 projects)
ProjectTest [210] Python, Java, 20 per language
JS
CLOVER [211] Python 845
TABLE XVI: Overview of Selected SE-Workflow Benchmarks.
@@
Code Synthesis & Understanding Methods2Test [212] Java 780,944
CRUXEval [213] Python 800
CRQBench [214] C++ 100
CriticBench [215] Python 3,825(across 5 tasks)
CodeScope [216] 8 PLs 13,390 (across 8 tasks)
CodeCriticBench [217] Various 1,517 (Easy), 1,084 (Medium),